# Credit Risk Meets Large Language Models: Building a Risk Indicator from Loan Descriptions in P2P Lending

### v1

@@
Abstract.
Peer-to-peer (P2P) lending has emerged as a distinctive financing mechanism, linking borrowers with lenders through online platforms. However, P2P lending faces the challenge of information asymmetry, as lenders often lack sufficient data to assess the creditworthiness of borrowers. This paper proposes a novel approach to address this issue by leveraging the textual descriptions provided by borrowers during the loan application process. Our methodology involves processing these textual descriptions using a Large Language Model (LLM), a powerful tool capable of discerning patterns and semantics within the text. Transfer learning is applied to adapt the LLM to the specific task at hand.
Our results derived from the analysis of the Lending Club dataset show that the risk score generated by BERT, a widely used LLM, significantly improves the performance of credit risk classifiers. However, the inherent opacity of LLM-based systems, coupled with uncertainties about potential biases, underscores critical considerations for regulatory frameworks and engenders trust-related concerns among end-users, opening new avenues for future research in the dynamic landscape of P2P lending and artificial intelligence.
@@ Abstract.
journal: TMISccs: Computing methodologies Information extractionccs: Applied computing Business process management systems
1. Introduction
@@
An inherent challenge in P2P lending is the presence of information asymmetry, wherein borrowers possess more and often superior information compared to lenders. To address this issue, platforms employ strategies to complement the conventional data provided in loan applications (Cummins et al., 2019). For instance, borrowers are frequently encouraged to provide a voluntary textual description describing the purpose of the loan and their particular situation. Despite the absence of formal verification, such voluntary disclosures have been observed to stimulate increased bidding activity among lenders. However, lenders may lack the expertise to effectively assess the creditworthiness of borrowers and may be influenced by different factors (Michels, 2012).
@@
Traditional credit scoring models do not harness the rich information embedded in the narratives submitted by loan applicants. Several approaches have tried to incorporate such information into credit risk models. The approaches include the extraction of linguistic metrics (Gao and Lin, 2015; Dorfleitner et al., 2016), the use of topic modeling to characterize the underlying themes (Xia et al., 2020; Zhang et al., 2020), or a combination of both (Siering, 2023).
In our work, we aim to use state-of-the-art Large Language Models (LLMs) to analyze textual descriptions and assign them a risk score. To achieve this, we will use BERT (Bidirectional Encoder Representations from Transformers) (Devlin et al., 2018), renowned for its outstanding capabilities in comprehending contextual nuances and semantic intricacies within textual data. BERT has demonstrated success in fine-tuning for classification tasks (Sun et al., 2019) and adapting to specific domains, such as biomedicine (Lee et al., 2019), as well as specialized tasks like spam classification (Tida and Hsu, 2022).
We will apply transfer learning to train BERT in distinguishing between default and non-default loans based on their textual descriptions. Subsequently, we will explore the viability of using the BERT output as a risk score and analyze how it relates to other variables. We integrate the BERT score into a credit risk classifier and show that it improves its predictive performance. Our results open new avenues for research by emphasizing the potential of LLMs to enhance credit risk models in P2P lending. Moreover, they underscore the need for greater transparency to streamline the adoption of these approaches, fostering trust among users and regulatory entities.
@@
This paper is organized as follows: Section 2 provides a comprehensive review of related work in the field of credit risk assessment and natural language processing. Section 3 presents an overview of LLMs and the BERT model. Section 4 describes the dataset used, detailing the data preprocessing steps and conducting an in-depth exploratory data analysis. Section 5 outlines the methodology, model architecture, and training procedures employed in integrating BERT into the credit risk assessment framework. Section 6 provides an analysis of the risk score generated by the BERT description processing. Section 7 discusses the results of our experiments, highlighting the improvements achieved by incorporating BERT-based textual analysis. Finally, Section 8 concludes the paper by summarizing key findings, discussing implications, and suggesting avenues for future research in the intersection of NLP and credit risk assessment.
2. Related work
@@
In their comprehensive analysis of risk-return modeling within the P2P lending market (Ariza-Garzón et al., 2021), the authors identify a discernible trend towards including new sources and types of information to improve risk and profit management models in the P2P market. The sources are very diverse and include transactional data (Xu et al., 2022), the topology of the lending-borrowing network (Li et al., 2018), data from social networks (Xu et al., 2016), or, more recently, facial features (Qi et al., 2022). Among them, the authors identify as a predominant trend the inclusion of textual data taken from statements describing the purpose of the loan.
@@
In a pioneer work exploring the impact of textual factors on peer-to-peer lending (Herzenstein et al., 2011), the authors analyze P2P loans including manually annotated narrative aspects, such as trustworthiness, economic hardship, hard work, success, morality, and religiosity. These aspects were combined with demographic variables and loan characteristics. Their results highlight that narratives regarding trustworthiness have a significant influence on decision-makers, i.e. credit lenders, in terms of securing loan funding, and that some of these narratives also play a significant role in loan performance.
@@
However, most subsequent studies typically use text mining or artificial intelligence methods to extract linguistic features or loan description topics. Regarding the use of linguistic features, the authors in (Gao and Lin, 2015) use machine learning and text mining techniques to quantify and extract linguistic features (e.g., readability, positivity, objectivity, and deception cues), and then build both explanatory econometric models and predictive models using such features. They find that they can indeed reflect borrowers’ creditworthiness and predict loan default. They also use a panel of investors and confirm that investors indeed value texts written by borrowers, but that they can also be deceived by some of the deception cues well established in the literature. Similarly, in (Dorfleitner et al., 2016), the authors include linguistic factors and the presence of social and emotional keywords and evaluate their impact on two European platforms. They found that text-derived variables influence the probability of funding, but not the probability of default. In (Wang et al., 2016), linguistic statistical features and abstract text features (including deception, subjectivity, sentiment, readability, personality, and mindset) are used to characterize text descriptions. They compare the performance of different classifiers based on the textual features and conclude that their performance is close to that of the classifiers using traditional financial features, but that adding textual features can improve the performance of the whole credit risk evaluation system.
@@
In topic modeling, the Latent Dirichlet Allocation model (LDA) has been widely used. In (Jiang et al., 2017), the authors use LDA to extract six topics from the loan descriptions whose meanings are obvious: assets, income and expenses, work, family, business, and agriculture. They also consider the number of characters in the descriptive text. They conclude that soft (qualitative) information can improve the performance of loan default prediction compared to existing methods based only on hard (quantitative) information and that soft features have a significant ability to discriminate loan defaults. Similarly, in (Yao et al., 2018) an LDA topic model is used to classify the loan titles into six purposes. The results show that the purpose significantly affects a borrower’s likelihood of obtaining financing. In particular, ambiguous titles, where a borrower does not adequately explain the purpose of the loan, significantly reduce the likelihood of successfully securing the loan. In (Xia et al., 2020), Xia et al. used a keyword clustering algorithm to extract the topics automatically. The algorithm uses keyword extraction based on term frequency-inverse document frequency (TF-IDF) and word embeddings using the Word2Vec model (Mikolov et al., 2013) (a neural network). Their results on three real-world datasets show that topic variables significantly improve predictability compared to using traditional information alone.
@@
More recently, Siering (Siering, 2023) investigated the effect of both aspects: topics and linguistic features. To extract topics, the author utilized a financial text analysis procedure (Loughran and McDonald, 2011) to build a domain dictionary. The identified topics describe the purpose of the loan, as well as the borrowers’ request for help, reliability, or appreciation. These topics were operationalized as binary indicator variables. The author employs text mining to create variables that measure aspects such as polarity, active orientation, readability, average sentence length, and word count. These variables are then used as input in a logistic regression. The results indicate that both linguistic and content-based factors aid in predicting a loan’s probability of default, with content-based factors being more significant than linguistic variables. The author also analyzes how the variables contribute to increasing or reducing the probability of default, and finds, for example, that expressing reliability has a positive effect.
@@
Yet, a significant gap becomes apparent in the application of state-of-the-art natural language techniques, including deep learning methods. In their work (Zhang et al., 2020), Zhang et al. study the transformer encoder’s ability to capture textual features from loan descriptions. These characteristics, alongside the traditional hard features derived from loan applications, are inputted into a neural network to predict the likelihood of loan default. The study highlights the effectiveness of transformers, showing better results when including textual loan descriptions compared to models that do not consider them.
In this work, we propose that the potential improvement could be further enhanced by using encoder-based LLMs, such as BERT, which has been successful in improving classification in other fields.
3. An overview of Large Language Models and BERT
@@
Large Language Models (LLMs) are built upon the Transformer architecture (Vaswani et al., 2017), leveraging attention mechanisms to enhance language comprehension. LLMs can be broadly categorized into three primary families, each distinguished by its architecture:
@@
  • Encoder-only, widely employed for language comprehension tasks such as text classification, named entity recognition, and extractive question answering. The most famous example is BERT (Devlin et al., 2018), and we will explain it in more detail below.
@@
  • Decoder-only, designed for generative tasks, exemplified by the well-known GPT models (Radford et al., 2019; Brown et al., 2020). It is employed in various tasks, including question answering (Pride et al., 2023), text summarization (Bhaskar et al., 2022), and programming code generation (Chen et al., 2021).
@@
  • Encoder-decoder models, suited for tasks demanding both language understanding and generation, such as language translation or text summarization. The most influential models are BART (Lewis et al., 2019) and T5 (Raffel et al., 2020).
@@
We will focus on BERT (Bidirectional Encoder Representations from Transformer) which is a Transformer-based language model introduced by Google researchers in 2018 (Devlin et al., 2018). Its architecture consists of a stack of encoders from the Transformer model. The bidirectional nature of BERT is key, as it considers both the left and right context of each word, enhancing its ability to understand context-dependent meanings and to be effective in language understanding tasks. Numerous studies have consistently shown that BERT is the most effective linguistic model for various of these tasks (Kriebel and Stitz, 2022; Stevenson et al., 2021). In contrast, the widely recognized GPT (Radford et al., 2019; Brown et al., 2020) model is built with a Transformer decoder stack, making it more suitable for language generation tasks.
@@
BERT stands as a milestone whose success has spurred the development of a diverse family of models that build upon its architecture. Some versions aim to achieve similar performance while having a smaller number of parameters, such as DistilBERT (a distilled version of BERT) (Sanh et al., 2019), or ALBERT (A Little BERT) (Lan et al., 2019). Others are adaptations to other languages such as CamemBERT (Martin et al., 2020) to French or BETO to Spanish (Cañete et al., 2020). Other proposals aimed to improve upon BERT by modifying some design decisions when pretraining BERT and also training the model longer, as in the case of RoBERTa (Robustly optimized BERT approach) (Liu et al., 2019), which resulted in improved contextualized representations and enhanced language understanding.
BERT has also been fine-tuned for specific tasks using transfer learning. For example, BioBERT (Lee et al., 2019) is a BERT model trained on biomedical documents to perform text mining tasks such as named entity recognition, relation extraction, and question answering. Other works fine-tune BERT to excel on text classification in well-known datasets (Sun et al., 2019) and adapted to specific applications on sentiment analysis (Gao et al., 2019) or spam classification (Tida and Hsu, 2022).
@@
In our research, we aim to develop a credit risk model by leveraging a BERT-based model. Specifically, our model will utilize the textual descriptions of loans provided by borrowers in their application forms, producing a predictive score indicative of the likelihood of loan default. We will assess the effectiveness of the score in a granting model using the well-established dataset from the Lending Club P2P lending platform.
4. Description of the dataset
We have used a public data set of the P2P lending company Lending Club [^1] https://www.kaggle.com/wordsforthewise/lending-club, which is widely used in credit risk publications and the most widely used when dealing with the P2P market (Ariza-Garzón et al., 2021).
The dataset includes loans granted by the company between 2007 and 2018, with a total number of 2,260,699 instances. However, we only consider loans whose final status is known, that is, that were either fully repaid or defaulted. Thus, we eliminate loans in transitory states (in a grace period, late, etc.) and exclude loans with no information on income and indebtedness. This led us to a dataset containing 1,347,681 instances.
Furthermore, we only use the loans that include the text description (desc variable), which is crucial for our study. For unknown reasons, we noticed that the descriptions were no longer available as of March 2014, so we further filtered the dataset considering only the loans from the beginning (June 2007) until that month.
After the desc variable cleaning, our dataset consists of 119,101 obligations, all of them including a description and with a target binary variable called Default. The target variable suffers the usual class imbalance problem (only 15.27% of loans defaulted), which is essential to take into account in the design of the experiments.
Table 1. Variable description.
Variable
Description

### v2

@@
Abstract
Peer-to-peer (P2P) lending connects borrowers with lenders through online platforms but faces significant information asymmetry, as lenders often lack sufficient data to assess the creditworthiness of borrowers. This paper addresses this challenge by leveraging BERT (Bidirectional Encoder Representations from Transformers), a Large Language Model (LLM) known for its ability to understand contextual nuances in text, to analyze borrowers’ loan descriptions.
We apply transfer learning to make BERT distinguish between default and non-default loans using the loan descriptions of the Lending Club dataset. The resulting BERT-generated risk score demonstrates solid competence in discriminating between default and non-default loans. Furthermore, integrating the BERT-generated risk score with traditional variables enhances classifier performance, which demonstrates the complementary nature of advanced language model outputs in refining credit risk assessment methodologies.
However, the opacity of LLMs and potential biases highlight the need for transparent regulatory frameworks. This study opens new research avenues in P2P lending and AI, emphasizing the importance of trust and transparency in adopting advanced credit risk models.
keywords:
@@ Abstract
[…]
[ucm] organization=Facultad de Informática, Universidad Complutense de Madrid, addressline=Calle del Prof. José García Santesmases, 9, city=Madrid, postcode=28040, state=Comunidad de Madrid, country=Spain
[…]
[itc] organization=Instituto de Tecnología del Conocimiento, Universidad Complutense de Madrid, addressline=Calle del Prof. José García Santesmases, 9, city=Madrid, postcode=28040, state=Comunidad de Madrid, country=Spain
{graphicalabstract}
{highlights}
Fine-tuned BERT successfully produces risk scores from loan descriptions.
Integrating BERT score with traditional variables boosts granting model performance.
The approach applies without manual annotation, reducing subjectivity and complexity.
Generating risk scores is instant with standard equipment.
Research is needed to enhance transparency and trust in these LLM-based approaches.
1 Introduction
@@
An inherent challenge in P2P lending is the presence of information asymmetry, wherein borrowers possess more and often superior information compared to lenders. To address this issue, platforms employ strategies to complement the conventional data provided in loan applications [1]. For instance, borrowers are frequently encouraged to provide a voluntary textual description describing the purpose of the loan and their particular situation. Despite the absence of formal verification, such voluntary disclosures have been observed to stimulate increased bidding activity among lenders. However, lenders may lack the expertise to assess the creditworthiness of borrowers effectively and may be influenced by different factors [2].
@@
Traditional credit scoring models do not harness the rich information embedded in the narratives submitted by loan applicants. Several approaches have tried to incorporate such information into credit risk models. The approaches include the extraction of linguistic metrics [3, 4], the use of topic modeling to characterize the underlying themes [5, 6], or a combination of both [7].
However, these methods have limitations in fully capturing the contextual nuances and semantic intricacies of loan descriptions. This study aims to address this gap by leveraging the capabilities of Large Language Models (LLMs), specifically BERT (Bidirectional Encoder Representations from Transformers) [8], which has shown exceptional performance in understanding and processing complex textual data.
BERT’s bidirectional training and ability to capture context at a granular level make it particularly suitable for tasks requiring deep semantic understanding. It has proven successful in fine-tuning for classification tasks [9], in particular domains such as biomedicine [10] and in specialized tasks like spam classification [11]. In this study, we apply transfer learning to train BERT to distinguish between default and non-default loans based on their textual descriptions.
Our work demonstrates BERT’s capability to effectively leverage descriptive data to generate a risk indicator that accurately classifies defaulted loans. Moreover, we show that integrating the BERT-generated risk score into traditional credit granting models can significantly enhance their predictive performance. This highlights BERT’s potential to improve the accuracy and reliability of credit risk models in P2P lending.
Furthermore, our findings highlight the need for greater transparency in adopting advanced models to build trust among users and regulatory entities. This study paves the way for future research, emphasizing the transformative potential of LLMs like BERT in the field of credit risk assessment.
@@
This paper is organized as follows: Section 2 provides a comprehensive review of related work in credit risk assessment and natural language processing. Section 3 presents an overview of LLMs and the BERT model. Section 4 describes the dataset used, detailing the data preprocessing steps and conducting an in-depth exploratory data analysis. Section 5 outlines the methodology, model architecture, and training procedures employed in integrating BERT into the credit risk assessment framework. Section 6 analyzes the risk score generated by the BERT description processing. Section 7 discusses the results of our experiments, highlighting the improvements achieved by incorporating BERT-based textual analysis. Finally, Section 8 concludes the paper by summarizing key findings, discussing implications, and suggesting avenues for future research in the intersection of NLP and credit risk assessment.
2 Related work
@@
In their comprehensive analysis of risk-return modeling within the P2P lending market [12], the authors identify a discernible trend towards including new sources and types of information to improve risk and profit management models in the P2P market. The sources are very diverse and include transactional data [13], the topology of the lending-borrowing network [14], data from social networks [15], or, more recently, facial features [16]. Among them, the authors identify as a predominant trend the inclusion of textual data taken from statements describing the purpose of the loan.
@@
In a pioneer work exploring the impact of textual factors on peer-to-peer lending [17], the authors analyze P2P loans including manually annotated narrative aspects, such as trustworthiness, economic hardship, hard work, success, morality, and religiosity. These aspects were combined with demographic variables and loan characteristics. Their results highlight that narratives regarding trustworthiness strongly influence decision-makers, particularly credit lenders, in their loan approval process. Additionally, some of these narratives play a substantial role in subsequent loan performance.
@@
However, most subsequent studies typically use text mining or artificial intelligence methods to extract linguistic features or loan description topics. Regarding the use of linguistic features, the authors in [3] use machine learning and text mining techniques to quantify and extract linguistic features (e.g., readability, positivity, objectivity, and deception cues), and then build both explanatory econometric models and predictive models using such features. They find that they can indeed reflect borrowers’ creditworthiness and predict loan default. They also use a panel of investors and confirm that investors indeed value texts written by borrowers, but that they can also be deceived by some of the deception cues well established in the literature. Similarly, in [4], the authors include linguistic factors and the presence of social and emotional keywords and evaluate their impact on two European platforms. They found that text-derived variables influence the probability of funding, but not the probability of default. In [18], linguistic statistical features and abstract text features (including deception, subjectivity, sentiment, readability, personality, and mindset) are used to characterize text descriptions. They compare the performance of different classifiers based on the textual features and conclude that their performance is close to that of the classifiers using traditional financial features, but that adding textual features can improve the performance of the whole credit risk evaluation system.
@@
As for topic modeling, the Latent Dirichlet Allocation model (LDA) has been widely used. In [19], the authors use LDA to extract six topics from the loan descriptions whose meanings are obvious: assets, income and expenses, work, family, business, and agriculture. They also consider the number of characters in the descriptive text. They conclude that soft (qualitative) information can improve the performance of loan default prediction compared to existing methods based only on hard (quantitative) information and that soft features have a significant ability to discriminate loan defaults. Similarly, in [20], an LDA topic model is used to classify the loan titles into six purposes. Their findings reveal that the stated purpose significantly influences a borrower’s chances of securing financing. Notably, ambiguous titles—where borrowers fail to clearly articulate the loan’s purpose—substantially diminish the likelihood of loan approval. In [5], Xia et al. used a keyword clustering algorithm for automatic topic extraction. Their method combines keyword extraction based on term frequency-inverse document frequency (TF-IDF) with word embeddings generated by the Word2Vec neural network model [21]. Analysis of three real-world datasets demonstrated that incorporating these topic variables significantly enhanced predictive accuracy compared to relying solely on traditional information.
@@
More recently, Siering [7] investigated the effect of both aspects: topics and linguistic features. To extract topics, the author utilized a financial text analysis procedure [22] to build a domain dictionary. The identified topics describe the purpose of the loan, as well as the borrowers’ request for help, reliability, or appreciation. These topics were operationalized as binary indicator variables. The author employs text mining to create variables that measure aspects such as polarity, active orientation, readability, average sentence length, and word count. These variables are then used as input in a logistic regression. The results indicate that both linguistic and content-based factors contribute to predicting loan default probability, with the latter demonstrating greater significance. Analysis of variable contributions shows that certain factors increase or decrease default likelihood. Notably, expressions of reliability correlate positively with loan repayment.
@@
Yet, a significant gap becomes apparent in applying state-of-the-art natural language processing techniques, including deep learning methods. In their work [6], Zhang et al. studied the transformer encoder’s ability to capture textual features from loan descriptions. These characteristics, alongside the traditional hard features derived from loan applications, were inputted into a neural network to predict the likelihood of loan default. The study highlights the effectiveness of transformers, showing better results when including textual loan descriptions compared to models that do not consider them. In this work, we explore using encoder-based LLMs, such as BERT, which has successfully improved classification in other fields.
3 An overview of Large Language Models and BERT
@@
Large Language Models (LLMs) are built upon the Transformer architecture [23], leveraging attention mechanisms to enhance language comprehension. LLMs can be broadly categorized into three primary families, each distinguished by its architecture:
  • 1.
@@
    Encoder-only, widely employed for language comprehension tasks such as text classification, named entity recognition, and extractive question answering. The most famous example is BERT [8], which will be explained in more detail below.
  • 2.
@@
    Decoder-only, designed for generative tasks, exemplified by the well-known GPT models [24, 25]. It is employed in various tasks, including question answering [26], text summarization [27], and programming code generation [28].
  • 3.
@@
    Encoder-decoder models, suited for tasks demanding both language understanding and generation, such as language translation or text summarization. The most influential models are BART [29] and T5 [30].
@@
We will focus on BERT (Bidirectional Encoder Representations from Transformer), which is a Transformer-based language model introduced by Google researchers in 2018 [8]. BERT’s architecture consists of a stack of encoders from the Transformer model. The bidirectional nature of BERT is key, as it considers both the left and right context of each word, enhancing its ability to understand context-dependent meanings and to be effective in language understanding tasks. Numerous studies have consistently shown that BERT is the most effective linguistic model for various of these tasks [31, 32]. Notably, BERT has 340 million parameters, while the widely recognized GPT-3 model has 175 billion, making BERT 514 times smaller than GPT-3 [25]. Given this significant size difference, BERT can be operated on standard home equipment for model inference, which greatly simplifies its use in practical scenarios. In contrast, GPT, built with a Transformer decoder stack, not only demands much more powerful equipment but is suited for language generation tasks.
@@
BERT stands as a milestone whose success has spurred the development of a diverse family of models that build upon its architecture. Some versions aim to achieve similar performance while having a smaller number of parameters, such as DistilBERT (a distilled version of BERT) [33], or ALBERT (A Little BERT) [34]. Others are adaptations to other languages such as CamemBERT [35] to French or BETO to Spanish [36]. Other proposals aimed to improve upon BERT by modifying some design decisions when pretraining BERT and also training the model longer, as in the case of RoBERTa (Robustly optimized BERT approach) [37], which resulted in improved contextualized representations and enhanced language understanding.
To further elucidate the role of BERT in specialized applications, it is crucial to understand its capacity for transfer learning and fine-tuning. Transfer learning involves using a pre-trained model like BERT, which has initially learned general language patterns from a large corpus to a specific task or dataset. This technique allows us to take advantage of the rich linguistic representations without needing extensive computation from scratch. Fine-tuning involves adjusting the pre-trained model’s parameters to capture the nuances of the target task or application field by further training with new instances from the new context. For example, BioBERT is a BERT model fine-tuned for biomedical text mining tasks like named entity recognition and question answering [10]. Other adaptations have targeted text classification and sentiment analysis in specific datasets [9, 38].
@@
Our research aims to develop a credit risk model by leveraging a BERT-based model. Specifically, we will fine-tune BERT for producing a predictive score indicative of the likelihood of loan default using the textual descriptions of loans provided by borrowers in their application forms. We will assess the scoring effectiveness in a granting model using the well-established dataset from the Lending Club P2P lending platform.
4 Description of the dataset
We have used a public data set of the P2P lending company Lending Club [^1] https://www.kaggle.com/wordsforthewise/lending-club, which is widely used in credit risk publications and the most widely used when dealing with the P2P market [39, 12]. However, instead of using the original dataset, which includes 2,260,699 loans granted by the company between 2007 and 2018, we will use a version modified for proposing granting models [40], used in [41, 42]. Since granting models determine which loans will be fully repaid, its estimation needs loans whose final status is known, that is, that were either fully repaid or defaulted. Thus, the dataset excludes loans in transitory states (in a grace period, late, etc.) and loans with no information on income and indebtedness which is essential to compute the input variables, resulting in 1,347,681 instances.
Additionally, the original dataset contains variables detailing the loan’s lifecycle and other post-application aspects (e.g., the interest rate). In contrast, our version only includes variables available at the time of application, which are those utilized by granting models.
Loan descriptions were inconsistently available, appearing only for certain loans between April 2008 and March 2014. To accurately assess the impact of textual descriptions on default prediction, our analysis focuses solely on the 119,101 loans that include the desc variable. Kolmogorov-Smirnov and chi-square tests were applied to quantitative and categorical variables to assess potential bias from filtering. The lack of significant differences indicates that the filtered dataset is representative of the original dataset.
Table 1: Variable description.
Variable Description